July 27th 2022

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • Sebastian working on updating database to decode the VTP readout
      • VTP did not get the
    • Hodoscope
      • still taking data every night, adjusting NINO threshold
      • HV scan was one
      • new HV crate was sent from Glasgow
      • some TDC extension coming loose : need to be remounted - fixed
    • Beamline
      • planning to do unser calibration but David was sick
      • Paul away next week but will give instruction
      • halladaq8 - MAC on
    • Parity crate
      • not affected
    • GEM
      • everything plugged in, waiting for the crates to boot
      • reset switch
      • patch panel fiber installed
      • runlist
      • C++ and python script ready
      • need to run into simplifying the authorization process